19 February 2015: Thursday

LDAP:

  • One issue.  A new employee that matches an existing record except it might not.  But it does.  Merged and account created.

Backups:

  • CoM/IS clones completed.  Disk Clones still running.
  • Performed the second round of tests for SR 66027568 (DD Compression) and sent results off to EMC.  They are pretty much the same as yesterday’s numbers.  That is not so good.  This could indicate a NIC congestion issue is the right answer.  Further tests using NetWorker (penguin-backup -> sn5/sn9 to see that) are needed.
  • SR66570742 (AFTD % Used) – no response to Tuesday’s email – updating the SR directly, perhaps I will get a response that way?
  • SR64026734 (Persistent Names – Scan issue) – installed the updated dvdetect program and ran the test – sent EMC the output – still generates error messages.  /sigh.
  • Put together preliminary plan for system maintenance to be done on February 23.

Footprints:

  • Issue 342136 – Don reports that having applied the patch, it still does not work.  Reported to BMC how I had got the patch to work and asked for a new webex so we can figure out why it is not working now.
  • Forced a logo that was too large to be smaller (made it look ugly, but it fits and does not block critical elements of the page now).

Email/Calendar:

  • Working on tying webauth into calendar.

Tags:

Leave a Reply

You must be logged in to post a comment.